# The second class particle in the half-line open TASEP

### v1

@@ 3.1. Second class particles in one shock
Let $M_{1},M_{2}$ be positive integers, and consider $\pi_{M_{1},M_{2}}$ be a signed permutation of the set $\{-M_{1}-M_{2}-1,\cdots,-1,1,\cdots,M_{1}+M_{2}+1\}$ such that $\pi_{M_{1},M_{2}}(i)=M_{1}+M_{2}+2-i$ and $\pi(-i)=-\pi(i)$ for $i=1,2,\cdots,M_{1}+M_{2}+1$. We assume the length $l(\pi_{M_{1},M_{2}})=m$ and let $\pi_{M_{1},M_{2}}=s_{k_{m}}s_{k_{m-1}}\cdots s_{k_{1}}$ be a minimal length dedomposition of $\pi_{M_{1},M_{2}}$ into the Coxeter generators (there can be many such decompositions, we only choose one of them).
@@
Because the dynamics on $\mathbb{N}$ and $-\mathbb{N}$ are symmetric, and the initial configurations (3.1) and (3.2) are filled with first class particles, therefore we can identity half-space TASEP $\eta_{t}$ (resp. half-space TASEP $\tilde{\eta}_{t}$) with initial condition (3.1) (resp. initial condition (3.2)) with the half-line open TASEP in Figure 1 Upper (resp. Figure 1 Bottom). That’s why we don’t change the notations. The reason is the same for the two-shock case in Section 3.4, we will not state it again later.
Both of above tricolor half-space TASEPs illustrate the second class particles in one shock. We aim to investigate the characteristics of the sole second-class particle in the half-space by examining the first process $\eta_{t}$. Simultaneously, we will explore the behavior of the multi-colored height functions in the half-space through the analysis of the second process $\tilde{\eta}_{t}$. Let $f(t)$ represent the position of the unique second class particle in the process $\eta_{t}$. Additionally, let $\mathbf{N}_{c}(x,t)$ denote the number of particles of color $c$ in the process $\tilde{\eta}_{t}$ at time $t$ that are located weakly to the right of $x\in\mathbb{R}_{\geq 0}$, where $c=1,2$, namely,
(3.3) $\mathbf{N}_{c}(x,t)=\sum_{z\geq x}\delta_{\tilde{\eta}_{t}(z),c},\quad\text{for}\ c=1,2.$
These quantities will be studied by establishing connections to a more straightforward process using the color-position symmetry.
@@ 3.3. Distribution identity
$\mathbb{P}((\iota\eta)_{t}^{gen;1}(M_{2}+1)<0)=\mathbb{P}(\eta_{t}^{gen;2}(M_{2}+1)<0).$
Considering the definition of $\eta_{t}^{gen;2}$ and the projection introduced in Section 3.2, the event that the color $\eta_{t}^{gen;2}(M_{2}+1)$ in the position $M_{2}+1$ of $\eta_{t}^{gen;2}$ is less than 0 is equivalent to the event in half-line TASEP (3.4) that the number of particles in the set $\{1,\cdots,M_{1}+M_{2}+1\}$ is at least $M_{1}+1$. Namely, we have the equation 3.6. ∎
@@
We use Coxeter generators data $s_{k_{1}},s_{k_{2}},\cdots,s_{k_{m}}$ introduced in Section 3.1 in the folloing. By the relation of $\tilde{\eta}_{t}$ and $\eta_{t}^{gen;1}$, we have
@@
Similar with the proof of Theorem 3.2, we can get 3.11. ∎
@@
(3.9) $\eta_{0}^{step(y)}(x)$ =$=\begin{cases}1,&\text{ if }x\leq y,\\ +\infty,&\text{ if }x>y.\end{cases}$
@@
(3.10) $\mathcal{N}^{y}(x,t)=\sum_{z\geq x}\delta_{\eta_{t}^{step(y)}(z),1}.$
@@
(3.11) $\left(\mathbf{N}_{1}(y,t),\mathbf{N}_{2}(y,t)\right)\stackrel{{\scriptstyle d}}{{=}}\\ \left(\mathcal{N}^{y}(1,t)-y,\min\left\{M_{1}+M_{2}+1-\mathcal{N}^{y}(1,t)+\mathcal{N}^{y}(M_{1}+M_{2}+2,t),M_{2}+1\right\}\right).$
Since we know very little about the asymptotic properties of $\mathcal{N}^{y}(x,t)$, this method is far to reach.
@@ 3.4. Second class particles in two shocks
We denote the position of the unique second class particle in the process $\xi_{t}$ by $g(t)$. Further, let $\mathsf{N}_{i}(x,t)$ be the number of particles of color $i$ in $\tilde{\xi}_{t}$ which are weakly to the right of $x\in\mathbb{R}$, where $i\in\{1,2,3\}$.
@@
Let $\pi_{1}$ be a signed permutation of the set $\{-2M-2N-1,\cdots,-M-1,M+1,\cdots,2M+2N+1\}$ such that $\pi_{1}(M+1+i)=\pi_{1}(2M+2N+1-i)$ and $\pi_{1}(-i)=-\pi_{1}(i)$, where $i=0,1,\cdots,M+2N$. And let $\pi_{2}$ be a signed permutation of the set $\{-M-N,\cdots,-1,1,\cdots,M+N\}$ such that $\pi_{2}(1+i)=\pi_{1}(M+N-i)$ and $\pi_{1}(-i)=-\pi_{1}(i)$, where $i=0,1,\cdots,M+N-1$. Consider a minimal length decomposition of permutation $\pi_{2}\cdot\pi_{1}$ into Coxeter generators: $\pi_{2}\cdot\pi_{1}=s_{\ell_{p}}s_{\ell_{p-1}}\cdots s_{\ell_{1}}$, where $l(\pi_{2}\cdot\pi_{1})=p$.
We use the data $s_{\ell_{1}},s_{\ell_{2}},\cdots,s_{\ell_{p}}$ to construct the first process $\eta_{t}^{gen;1}$ introduced in Section 2.3. We can get $\xi_{t}$ from $\eta_{t}^{gen;1}$(with Coxeter generators data $s_{\ell_{1}},s_{\ell_{2}},\cdots,s_{\ell_{p}}$) by identifying the particle with color $M+N+1$ as the second class particle (denote by 2), particles with colors less than $M+N+1$ as first class particles (denote by 1), and particles with colors greater than $M+N+1$ are holes (denote by +$\infty$+ ∞). That means:
@@
$\left(\mathsf{N}_{1}(1,t),\mathsf{N}_{2}(1,t),\mathsf{N}_{3}(1,t)\right)\stackrel{{\scriptstyle d}}{{=}}\left(\#\{i:i<0,\eta_{t}^{gen;2}(i)\geq 1\},\right.\\ \left.\#\{i:1\leq i\leq M,\eta_{t}^{gen;2}(i)\geq 1\},\#\{i:M+1\leq i\leq M+N+1,\eta_{t}^{gen;2}(i)\geq 1\}\right),$
where we use Coxeter generators data $s_{\ell_{1}},s_{\ell_{2}},\cdots,s_{\ell_{p}}$ to construct $\eta_{t}^{gen;2}$(recall the definition introduced in Section 2.3). Considering the definition of $\eta_{t}^{gen;2}$ and the projection introduced in Section 3.2, the color of the particle in position $M+N+1$ is less than 0 in $\eta_{t}^{gen;2}$ is equivalent to the event in half-line TASEP 3.4 that the number of particles in the set $\{M+1,\cdots,2M+2N+1\}$ is at least $M+N+1$. Namely, we have the equation 3.14. Similarly, we have the following distribution identity:
#$\#\{i:i<0,\eta_{t}^{gen;2}(i)\geq 1\}\ \stackrel{{\scriptstyle d}}{{=}}\ \#\{i:i>0,\eta_{t}^{gen;2}(i)\leq-1\}\ \stackrel{{\scriptstyle d}}{{=}}\ \mathcal{N}(1,t),\\ \#\{i:1\leq i\leq M,\eta_{t}^{gen;2}(i)\geq 1\}\ \stackrel{{\scriptstyle d}}{{=}}\ \min\{M,M+N-\mathcal{N}(1,t)+\mathcal{N}(M+N+1,t)\}\\ \#\{i:M+1\leq i\leq M+N+1,\eta_{t}^{gen;2}(i)\geq 1\}\stackrel{{\scriptstyle d}}{{=}}\min\bigl{\{}N+1,M+N+1-\mathcal{N}(M+N+1,t)\bigr{.}\\ \bigl{.}+\mathcal{N}(2M+2N+2,t)+\max\{0,N-\mathcal{N}(1,t)+\mathcal{N}(M+N+1,t)\}\bigr{\}},$
which lead to the proof of 3.15. ∎
@@ 4.1. Stationary measure of half-line open TASEP
According to Liggett’s results on half-line open ASEP [Ligg75], we can characterize the stationary measure of half-line TASEP with step initial condition (denoted by $\eta_{t}^{step}$) as follows:
@@
Let parameter $\alpha\in(0,1)$ corresponds to the density of particles that the reservoir imposes at site 1. When $\alpha\leq\frac{1}{2}$, the process $\eta_{t}^{step}$ admits stationary measures that are product i.i.d. Bernoulli with parameter $\alpha$. When $\alpha>\frac{1}{2}$, stationary measures (denoted by $\mu_{1/2}^{\alpha}$) of process $\eta_{t}^{step}$ are spatially correlated and more complicated: there is a rarefaction fan with a density of particles $\alpha$ near the origin and density $\frac{1}{2}$ at +$\infty$+ ∞.
Following the seminal work of [DEHP93], Grosskinsky [Gro04] has given a explicit characterisation of the stationary measure $\mu_{1/2}^{\alpha}$ of half-line open TASEP with a matrix product ansatz:
@@
The $L$ dimensional distributions of stationary measures $\mu_{1/2}^{\alpha}$ for the process $\eta_{t}^{step}$ are given by the following matrix product ansatz:
where $L\in\mathbb{N}$, $(\eta_{1},\cdots,\eta_{L})\in\{0,1\}^{L}$, the nonnegative matrices $D$, $E$ and vectors $\left\langle w\right|$ and $\left|v\right\rangle$ satisfy the following algebraic relations with $c=\frac{1}{4}$:
@@
(4.5) $\left\langle w\right|\left|v\right\rangle=1.$
@@
$=$ $(c/\alpha)^{\tau_{0}}\left\langle w\right|\underbrace{D\ldots D}_{\sigma_{1}}\underbrace{E\ldots E}_{\tau_{1}}\underbrace{D\ldots D}_{\sigma_{2}}\cdots\underbrace{E\ldots E}_{\tau_{n-1}}\underbrace{D\ldots D}_{\sigma_{n}}\underbrace{E\ldots E}_{\tau_{n}}\left|v\right\rangle$
::=:= $(c/\alpha)^{\tau_{0}}F(\sigma_{1},\tau_{1},\cdots,\sigma_{n},\tau_{n}),$
where we use the algebraic relation (4.3) and the expression (4.7) of $\eta$. Our main task is to express $F(\sigma_{1},\tau_{1},\cdots,\sigma_{n},\tau_{n})$ as a linear combination of $\left\langle w\right|D^{k}\left|v\right\rangle$.
@@
Now we are in a position to extend the definition of one-layer DEHP-tree to $n$-layer DEHP-tree, and then give the explicit combinatorial formula of $F(\sigma_{1},\tau_{1},\cdots,\sigma_{n},\tau_{n})$ with arbitrary $n$ pairs of blocks $\prod_{i=1}^{n}\underbrace{D\ldots D}_{\sigma_{i}}\underbrace{E\ldots E}_{\tau_{i}}$. We introduce the notation $\tilde{\Psi}_{i}=\sum_{j=1}^{i}\sigma_{j}$ and $\tilde{\Phi}_{i}=\sum_{j=1}^{i}\tau_{j}$.
@@
The $n$-layer DEHP-tree of $(\tilde{\Psi}_{n},\tilde{\Phi}_{n})$ is a rooted tree with DEHP weight, which is defined by the following recursive way:
  • start from an one-layer DEHP-tree of $(\tilde{\Psi}_{1},\tilde{\Phi}_{1})$;
@@
We can give the explicit formula of $F(\sigma_{1},\tau_{1},\cdots,\sigma_{n},\tau_{n})$ by the above $n$-layer DEHP-tree:
(4.9) $F(\sigma_{1},\tau_{1},\cdots,\sigma_{n},\tau_{n})=\sum_{k=0}^{\tilde{\Psi}_{n}}Z_{k}(\tilde{\Psi}_{n},\tilde{\Phi}_{n})\left\langle w\right|D^{k}\left|v\right\rangle.$
@@
We can conclude the proof from the definition 4.5. ∎
@@
For any $\eta$ with the form 4.7, the stationary measure 4.1 has the explicit formula:
(4.10) $\mu_{1/2}^{\alpha}\{\xi\in\{0,1\}^{\mathbb{N}}:\xi_{1}=\eta_{1},\cdots,\xi_{n}=\eta_{L}\}=\sum_{k=0}^{\tilde{\Psi}_{n}}\frac{c^{\tau_{0}}Z_{k}(\tilde{\Psi}_{n},\tilde{\Phi}_{n})}{2^{k}\alpha^{\tau_{0}}}\left(1+\frac{\alpha-\frac{1}{2}}{\alpha}k\right),$
where $c=\frac{1}{4}$ and $Z_{k}(\Psi_{n},\Phi_{n})$ is given by the partition function of $(k,0)$ in the $n$-layer DEHP-tree.
Remark 4.8.
One can check the construction of DEHP-tree in Definition 1.1 is equivalent to the above construction. The difference is that we incorporate the term $\left\langle w\right|\underbrace{E\ldots E}_{\tau_{0}}$ into the Definition 1.1. The reason we do this is that the expression of formula is more neat:
(4.11) $\mu_{1/2}^{\alpha}\{\xi\in\{0,1\}^{\mathbb{N}}:\xi_{1}=\eta_{1},\cdots,\xi_{n}=\eta_{L}\}=\sum_{k=0}^{\Psi_{1}(\eta)}\frac{Z_{k}(\eta)}{2^{k}}\cdot\left(1+\frac{\alpha-\frac{1}{2}}{\alpha}k\right).$
where $\Psi_{1}(\eta)$ and $Z_{k}(\eta)$ are defined in Section 1.3.

### v2

@@ 3.1. Second class particles in one shock
Let $M_{1},M_{2}$ be positive integers, and consider $\pi_{M_{1},M_{2}}$ be a signed permutation of the set $\{-M_{1}-M_{2}-1,\cdots,-1,1,\cdots,M_{1}+M_{2}+1\}$ such that $\pi_{M_{1},M_{2}}(i)=M_{1}+M_{2}+2-i$ and $\pi(-i)=-\pi(i)$ for $i=1,2,\cdots,M_{1}+M_{2}+1$. We assume the length $l(\pi_{M_{1},M_{2}})=m$ and let $\pi_{M_{1},M_{2}}=s_{k_{m}}s_{k_{m-1}}\cdots s_{k_{1}}$ be a minimal length decomposition of $\pi_{M_{1},M_{2}}$ into the Coxeter generators (there can be many such decompositions, we only choose one of them).
@@
In this paper, we consider several projections of fully colored half-space TASEP where we map colors $(-\infty,K_{1}]$ to 1, colors $[K_{1}+1,K_{2}]$ to 2, $\ldots$, colors $[K_{N-1}+1,K_{N}]$ to $N$, and colors $[K_{N},+\infty]$ to +$\infty$+ ∞, for some fixed $N\in\mathbb{N}$ and $-\infty<K_{1}<K_{2}<\cdots<K_{N}<+\infty$. Note that the set of colors is linearly ordered, and particle with a smaller color interacts with a particle with a larger color as a particle with hole. It follows immediately from verifying the marginal transition rates that these projections have the same laws as the multi-species extensions of the half-line open TASEP.
Remark 3.2.
Because the dynamics on $\mathbb{N}$ and $-\mathbb{N}$ are symmetric, and the initial configurations (3.1) and (3.2) are filled with first class particles, therefore we can identity half-space TASEP $\eta_{t}$ (resp. half-space TASEP $\tilde{\eta}_{t}$) with initial condition (3.1) (resp. initial condition (3.2)) with the half-line open TASEP in Figure 1 Upper (resp. Figure 1 Bottom). That’s why we don’t change the notations. The reason is the same for the two-shock case in Section 3.4.
@@ 3.3. Distribution identity
$\mathbb{P}((\iota\eta)_{t}^{gen;1}(M_{2}+1)<0)=\mathbb{P}(\eta_{t}^{gen;2}(M_{2}+1)<0),$
where $\eta_{t}^{gen;2}(M_{2}+1)<0$ means the color in the position $M_{2}+1$ is less than 00 in the random configuration $\eta_{t}^{gen;2}$. Recall that we get the random configuration $\eta_{t}^{gen;2}$ by the following way: we start from the identity configuration, and then we perform the continuous time process for time $t$, finally we sort the colors on the interval $[1,M_{1}+M_{2}+1]$ in a decreasing order, at the same time we sort the colors on the interval $[-M_{1}-M_{2}-1,-1]$ in a decreasing order.
Consider the projection introduced in Section 3.2: we say that if a color is $\leq-1$, then we have a particle; while if a color is $\geq 1$, then we have a hole. Under this projection, we can interpret $\eta_{t}^{gen;2}$ as the standard half-line open TASEP: at time $t=0$, we have the step initial condition; after running the dynamics for time $t$, we rearrange the holes on the interval $[1,M_{1}+M_{2}+1]$ to the left and the particles to the right. Thus, the event $\eta_{t}^{gen;2}(M_{2}+1)<0$ is equivalent to the event in half-line TASEP (3.3) that the number of particles in the set $\{1,\cdots,M_{1}+M_{2}+1\}$ is at least $M_{1}+1$. Namely, we have the equation 3.5. ∎
@@
We use Coxeter generators data $s_{k_{1}},s_{k_{2}},\cdots,s_{k_{m}}$ introduced in Section 3.1 in the following. By the relation of $\tilde{\eta}_{t}$ and $\eta_{t}^{gen;1}$, we have
@@
Similar with the proof of Theorem 3.3, we can get 3.7. ∎
@@
$\eta_{0}^{step(y)}(x)$ =$=\begin{cases}1,&\text{ if }x\leq y,\\ +\infty,&\text{ if }x>y.\end{cases}$
@@
$\mathcal{N}^{y}(x,t)=\sum_{z\geq x}\delta_{\eta_{t}^{step(y)}(z),1}.$
@@
$\left(\mathbf{N}_{1}(y,t),\mathbf{N}_{2}(y,t)\right)\stackrel{{\scriptstyle d}}{{=}}\\ \left(\mathcal{N}^{y}(1,t)-y,\min\left\{M_{1}+M_{2}+1-\mathcal{N}^{y}(1,t)+\mathcal{N}^{y}(M_{1}+M_{2}+2,t),M_{2}+1\right\}\right).$
We will address the asymptotic analysis in our future work.
@@ 3.4. Second class particles in two shocks
For the same reason as in Remark 3.2, the model $\xi_{t}$ is equivalent to the half-line open TASEP shown in the upper part of Figure 3, and the model $\tilde{\xi}_{t}$ is equivalent to the one shown in the bottom part of Figure 3. We use the same notations $g(t)$ and $\mathsf{N}_{i}(x,t)$ as defined in Section 1.2.
@@
Let $\pi_{1}$ be a signed permutation of the set $\{-2M-2N-1,\cdots,-M-1,M+1,\cdots,2M+2N+1\}$ such that $\pi_{1}(M+1+i)=2M+2N+1-i$ and $\pi_{1}(-i)=-\pi_{1}(i)$, where $i=0,1,\cdots,M+2N$. And let $\pi_{2}$ be a signed permutation of the set $\{-M-N,\cdots,-1,1,\cdots,M+N\}$ such that $\pi_{2}(1+i)=M+N-i$ and $\pi_{1}(-i)=-\pi_{1}(i)$, where $i=0,1,\cdots,M+N-1$. Consider a minimal length decomposition of permutation $\pi_{2}\cdot\pi_{1}$ into Coxeter generators: $\pi_{2}\cdot\pi_{1}=s_{\ell_{p}}s_{\ell_{p-1}}\cdots s_{\ell_{1}}$, where $l(\pi_{2}\cdot\pi_{1})=p$.
We use the data $s_{\ell_{1}},s_{\ell_{2}},\cdots,s_{\ell_{p}}$ to construct the first process $\eta_{t}^{gen;1}$ introduced in Section 2.3. We can get $\xi_{t}$ from $\eta_{t}^{gen;1}$(with Coxeter generators data $s_{\ell_{1}},s_{\ell_{2}},\cdots,s_{\ell_{p}}$) by identifying the particle with color $M+N+1$ as the second class particle (denote by 2), particles with colors less than $M+N+1$ as first class particles (denote by 1), and particles with colors greater than $M+N+1$ are holes (denote by +$\infty$+ ∞). This implies:
@@
(3.12) $\left(\mathsf{N}_{1}(1,t),\mathsf{N}_{2}(1,t),\mathsf{N}_{3}(1,t)\right)\stackrel{{\scriptstyle d}}{{=}}\left(\#\{i:i<0,\eta_{t}^{gen;2}(i)\geq 1\},\right.\\ \left.\#\{i:1\leq i\leq M,\eta_{t}^{gen;2}(i)\geq 1\},\#\{i:M+1\leq i\leq M+N+1,\eta_{t}^{gen;2}(i)\geq 1\}\right),$
where we use Coxeter generators data $s_{\ell_{1}},s_{\ell_{2}},\cdots,s_{\ell_{p}}$ to construct $\eta_{t}^{gen;2}$. Recall that we get the random configuration $\eta_{t}^{gen;2}$ by the following way: we start from the identity configuration; and then we perform the continuous time process for time $t$; after that we separately sort the colors on the interval $[1,M+N]$ and $[-M-N,-1]$ in a decreasing order at the same time; finally we separately sort the colors on the interval $[M+1,2M+2N+1]$ and $[-2M-2N-1,-M-1]$ in a decreasing order.
Consider the projection introduced in Section 3.2: we say that if a color is $\leq-1$, then we have a particle; while if a color is $\geq 1$, then we have a hole. Under this projection, we can interpret $\eta_{t}^{gen;2}$ as the standard half-line open TASEP: at time $t=0$, we have the step initial condition; after running the dynamics for time $t$, we rearrange the holes on the interval $[1,M+N]$ to the left and the particles to the right; finally, we rearrange the holes on the interval $[M+1,2M+2N+1]$ to the left and the particles to the right. Thus, the event $\eta_{t}^{gen;2}(M+N+1)<0$ is equivalent to the event in half-line TASEP (3.3) that the number of particles $\mathcal{N}(N+M+1,t)-\mathcal{N}(2N+2M+2,t)+\min\{N,\mathcal{N}(1,t)-\mathcal{N}(N+M+1,t)\}$ are at least $M+N+1$. Namely, we have the equation 3.10.
Under the above projection, we can also establish the following identities between the events in $\eta_{t}^{gen;2}$ and the events in standard half-line open TASEP:
#$\#\{i:i<0,\eta_{t}^{gen;2}(i)\geq 1\}=\#\{i:i>0,\eta_{t}^{gen;2}(i)\leq-1\}\ =\mathcal{N}(1,t),\\ \#\{i:1\leq i\leq M,\eta_{t}^{gen;2}(i)\geq 1\}=\min\{M,M+N-\mathcal{N}(1,t)+\mathcal{N}(M+N+1,t)\}\\ \#\{i:M+1\leq i\leq M+N+1,\eta_{t}^{gen;2}(i)\geq 1\}=\min\bigl{\{}N+1,M+N+1-\mathcal{N}(M+N+1,t)\bigr{.}\\ \bigl{.}+\mathcal{N}(2M+2N+2,t)+\max\{0,N-\mathcal{N}(1,t)+\mathcal{N}(M+N+1,t)\}\bigr{\}},$
which lead to the proof of 3.11 by using the joint distribution identity (3.12). ∎
@@ 4.1. Stationary measure of half-line open TASEP
Let the parameter $\alpha\in(0,1)$ represent the particle density imposed by the reservoir at site 1. When $\alpha\leq\frac{1}{2}$, Liggett’s result [Ligg75, Theorem 1.8] characterizes the stationary distribution of the half-line open TASEP with step initial condition (denoted by $\eta_{t}^{step}$) as follows:
@@
For $\alpha\leq\frac{1}{2}$, the half-line open TASEP with step initial condition $\eta_{t}^{\mathrm{step}}$ converges to a product measure of i.i.d. Bernoulli random variables with parameter $\alpha$.
When $\alpha>\frac{1}{2}$, Liggett [Ligg75] showed that the limiting stationary measures of the process $\eta_{t}^{step}$ are spatially correlated and have an asymptotic density of $\frac{1}{2}$. Based on physical arguments, these limiting stationary measures are conjectured to coincide with the stationary measure described in [Gro04, Theorem 3.2] with stationary current $c=\frac{1}{4}$. However, establishing this connection rigorously remains an open and technically challenging problem. In this paper, we present several theorems that are conditional on this conjecture. For completeness, we state below the stationary measure given in [Gro04, Theorem 3.2] for the case $c=\frac{1}{4}$ and the above conjecture.
@@
Let $D$, $E$ be non-negative matrices, and $\left\langle w\right|$, $\left|v\right\rangle$ be vectors satisfying the following algebraic relations with $c=\frac{1}{4}$:
@@
(4.4) $\langle w|v\rangle=1.$
Then for $\alpha>\frac{1}{2}$, the probability measure $\mu_{1/2}^{\alpha}$ defined via the following matrix product ansatz:
for any $L\in\mathbb{N}$ and $(\eta_{1},\cdots,\eta_{L})\in\{0,1\}^{L}$, is stationary for the dynamics of half-line open TASEP.
Conjecture 4.3.
For $\alpha>\frac{1}{2}$, the half-line open TASEP with step initial condition $\eta_{t}^{\mathrm{step}}$ converges to the probability measure $\mu_{1/2}^{\alpha}$ defined in 4.5.
The main result of this section is the following Theorem.
Theorem 4.4.
For any $L\in\mathbb{N}$ and any state $\eta=(\eta_{1},\cdots,\eta_{L})\in\{0,1\}^{L}$, we define the DEHP-tree of $\eta$ by definition 1.1. The stationary measure 4.5 has the explicit combinatorial formula:
(4.6) $\mu_{1/2}^{\alpha}\{\xi\in\{0,1\}^{\mathbb{N}}:\xi_{1}=\eta_{1},\cdots,\xi_{n}=\eta_{L}\}=\sum_{k=0}^{\Psi_{1}(\eta)}\frac{Z_{k}(\eta)}{2^{k}}\cdot\left(1+\frac{\alpha-\frac{1}{2}}{\alpha}k\right),$
where $\Psi_{1}(\eta)$ and $Z_{k}(\eta)$ are defined in Section 1.3.
In the rest of this section, we will analyze the structure of matrix product ansatz (4.5) and give a proof of theorem 4.4.
@@
::=:= $F(\tau_{0},\sigma_{1},\tau_{1},\cdots,\sigma_{n},\tau_{n}),$
Our main task is to express $F(\tau_{0},\sigma_{1},\tau_{1},\cdots,\sigma_{n},\tau_{n})$ as a linear combination of $\left\langle w\right|D^{k}\left|v\right\rangle$.
@@
Now we are in a position to extend the definition of one-layer DEHP-tree to multi-layer DEHP-tree, and then give the explicit combinatorial formula of $F(\tau_{0},\sigma_{1},\tau_{1},\cdots,\sigma_{n},\tau_{n})$ with arbitrary pairs of blocks $\prod_{i=0}^{n}\underbrace{D\ldots D}_{\sigma_{i}}\underbrace{E\ldots E}_{\tau_{i}}$, where $\sigma_{0}=0$. We introduce the notation $\tilde{\Psi}_{i}=\sum_{j=0}^{i}\sigma_{j}$ and $\tilde{\Phi}_{i}=\sum_{j=0}^{i}\tau_{j}$.
@@
The multi-layer DEHP-tree of $(\tilde{\Psi}_{n},\tilde{\Phi}_{n})$ is a rooted tree with DEHP weight, which is defined by the following recursive way:
  • start from an one-layer DEHP-tree of $(\tilde{\Psi}_{0},\tilde{\Phi}_{0})$;
@@
We can give the explicit formula of $F(\tau_{0},\sigma_{1},\tau_{1},\cdots,\sigma_{n},\tau_{n})$ by the above multi-layer DEHP-tree:
(4.10) $F(\tau_{0},\sigma_{1},\tau_{1},\cdots,\sigma_{n},\tau_{n})=\sum_{k=0}^{\tilde{\Psi}_{n}}Z_{k}(\tilde{\Psi}_{n},\tilde{\Phi}_{n})\left\langle w\right|D^{k}\left|v\right\rangle.$
@@
By definition, we have $F(\tau_{0},\sigma_{1},\tau_{1},\cdots,\sigma_{n},\tau_{n})=\left\langle w\right|\prod_{j=0}^{n}\underbrace{D\ldots D}_{\sigma_{j}}\underbrace{E\ldots E}_{\tau_{j}}\left|v\right\rangle$, where $\sigma_{0}=0$. We calculate the product step by step. First, $\left\langle w\right|\underbrace{E\ldots E}_{\tau_{0}}$ is given by the one-layer DEHP-tree: $Z_{0}(0,\tau_{0})\left\langle w\right|$. If we already have $\left\langle w\right|\prod_{j=0}^{i-1}\underbrace{D\ldots D}_{\sigma_{j}}\underbrace{E\ldots E}_{\tau_{j}}=\sum_{k=0}^{\tilde{\Psi}_{i-1}}Z_{k}(\tilde{\Psi}_{i-1},\tilde{\Phi}_{i-1})\left\langle w\right|D^{k}$, then $\left\langle w\right|\prod_{j=1}^{i}\underbrace{D\ldots D}_{\sigma_{j}}\underbrace{E\ldots E}_{\tau_{j}}$ can be expressed by
$\sum_{k=0}^{\tilde{\Psi}_{i-1}}Z_{k}(\tilde{\Psi}_{i-1},\tilde{\Phi}_{i-1})\left\langle w\right|D^{k}\underbrace{D\ldots D}_{\sigma_{i}}\underbrace{E\ldots E}_{\tau_{i}}.$
By the proof of Proposition 4.6 and the definition of $i$-layer DEHP-tree, the above expression is equivalent to $\sum_{k=0}^{\tilde{\Psi}_{i}}Z_{k}(\tilde{\Psi}_{i},\tilde{\Phi}_{i})\left\langle w\right|D^{k}$. Finally, at step $n$, we can get the result (4.10). ∎
@@
For any $\eta$ with the form 4.8, the stationary measure 4.5 has the explicit formula:
(4.11) $\mu_{1/2}^{\alpha}\{\xi\in\{0,1\}^{\mathbb{N}}:\xi_{1}=\eta_{1},\cdots,\xi_{n}=\eta_{L}\}=\sum_{k=0}^{\tilde{\Psi}_{n}}\frac{Z_{k}(\tilde{\Psi}_{n},\tilde{\Phi}_{n})}{2^{k}}\left(1+\frac{\alpha-\frac{1}{2}}{\alpha}k\right),$
where $Z_{k}(\Psi_{n},\Phi_{n})$ is given by the partition function of $(k,0)$ in the multi-layer DEHP-tree of $(\tilde{\Psi}_{n},\tilde{\Phi}_{n})$.
Now we can provide the proof of theorem 4.4.
Proof.
In the above, we use the recursive way to give the explicit formula of stationary measure 4.5. Now we use the notation in Section 1.3 to give a direct way. In definition 4.7, we construct the multi-layer DEHP-tree by starting from node $(\tilde{\Psi}_{0},\tilde{\Phi}_{0})$, which is exactly the starting node $(\Psi_{1}(\eta),\Phi_{0}(\eta))$ in definition 1.1. By the definition 4.5, any node in the multi-layer DEHP-tree 4.7 have either one child or two children. When $(x,y)\in\{(x,y):x=\Psi_{i}(\eta)\ \text{and}\ \Phi_{i}(\eta)<y\leq\Phi_{i-1}(\eta)\ \text{for}\ 1\leq i\leq A(\eta)+1\}$, the node $(x,y)$ has only one child. When $(x,y)\in\{(x,y):x=\Psi_{i}(\eta)\ \text{and}\ 1\leq y\leq\Phi_{i}(\eta)\ \text{for}\ 1\leq i\leq A(\eta)\}\cup\{(x,y):\Psi_{i+1}(\eta)<x<\Psi_{i}(\eta),1\leq y\leq\Phi_{i}(\eta)\ \text{for}\ 1\leq i\leq A(\eta)\}$, the node $(x,y)$ has two children. One can easily find that the above observations are exactly the rules in definition 1.1. The proof is done. ∎